OpenStudy (anonymous):

(x+3y)(x^2+7xy-9y^2)

OpenStudy (anonymous):

Distributive property. What do you need to know?

OpenStudy (anonymous):

\[(x+3y)(x^2+7xy-9y^2)=x(x^2+7xy-9y^2)+3y(x^2+7xy-9y^2)\]\[=x^3+7x^2y-9xy^2-3yx^2+21xy^2-27y^3=x^3+4x^2y+12xy^2-27y^3\]